# Rapid Word Learning Through Meta In-Context Learning

### v2

@@ Appendix A Word Usage Dataset Creation
Since the basic units of our focus are words (as opposed to word pieces in other tokenization schemes), we need to identify words in the text. To achieve this, we apply the same word-level tokenization to all datasets (for consistency) and mark word boundaries by whitespace during preprocessing. Models trained from scratch use this word-level tokenization. When the text is used in finetuning Llama-3, which comes with its pre-trained subword tokenizer, we remove the unnatural spaces introduced by the word-level tokenization and tokenize the text again with Llama-3 tokenizer, so the text format becomes closer to its pre-training data (See the Finetuning paragraph in Appendix B for further details of this process). For CHILDES data, we preprocess the data in the same way as Yedetore et al. (2023) did, which uses children’s input in the North American English portion, [^13] The version of CHILDES data we use is different from that of Yedetore et al. (2023), and the current version on the official webpage https://childes.talkbank.org/access/Eng-NA/ has also changed from our version. but we do not split and unk the data at the preprocessing stage. For BabyLM data, we use the data in the 10M track of the BabyLM Challenge 2023, which mixes 10 portions, each from a different data source (child- or adult-oriented, speech transcription or written text like Wikipedia). We exclude the QED portion for its poor quality (also mentioned in the 2nd BabyLM Challenge). We apply word-level tokenization on untokenized portions, and then split the text into sentences using heuristics. We use spaCy for all word-level tokenization along with Part-Of-Speech tagging. We lowercase all text before preprocessing to unify the capitalization of words in different places. We deduplicate sentences and remove sentences having less than 1 word (not counting punctuation).
@@ Appendix B Model and Training Configurations
We slightly modify the configuration of Pythia-160M (Biderman et al., 2023), which uses the Transformer architecture GPT-NeoX (Andonian et al., 2023). The configuration has $12$ layers and a hidden dimension size of $768$. We change the vocabulary size according to the corresponding dataset, as shown in Table 5. We also include three special tokens in the vocabulary: the placeholder token [new-token], the separator token <sep>, and <unk>, as mentioned in Section 4. We change the Pythia configuration to tie the input and output embeddings. This makes the model parameter counts smaller, $86.7$M and $102.5$M for the model trained on CHILDES and BabyLM-10M, respectively. For both models, we use batch size (i.e., number of episodes/sequences per batch) $8$ and AdamW optimizer (Loshchilov and Hutter, 2019) with initial learning rate $3\times 10^{-4}$, and reduce the learning rate by multiplying $0.1$ when the validation loss has stopped improving for $2$ epochs. We apply weight decay $0.07$ and $0.15$ when training on the CHILDES and BabyLM-10M datasets, respectively. Other configurations, such as no dropout, are kept the same as Pythia-160M. For each setting, we run $3$ times with random seed $\{0,1,2\}$. Each run is performed on a single V100 GPU for 30 epochs (9–18 hours).
@@
We finetune Llama-3 8B (Meta AI, 2024) with Minnow on each of the CHILDES and BabyLM-10M datasets, but we refer to the models finetuned on BabyLM-10M by default, as we mentioned in Section 5. We finetune from both the pre-trained and instruction-tuned variants of Llama-3 8B, but we refer to the models finetuned from the pre-trained variant by default, presenting results of finetuning from the instruction-tuned variant only in the generative settings, where their performance may differ significantly due to their different capabilities to follow the prompt. We use two reserved special tokens in Llama-3 tokenizer vocabulary as the placeholder token and the separator token. To make the tokenization more natural to the model’s pre-training data, we clean up tokenization spaces in the text (e.g., the space before “,”, “.”, or “’s”) introduced by the word-level tokenization during preprocessing and make the placeholder token absorbs any preceding spaces of the word. Finetuning is minimally parameter-efficient: We finetune only the input and output embeddings of the two special tokens, while freezing all other parameters. Before finetuning, the input/output embedding of either token is initialized to the mean of all input/output embeddings (Hewitt, 2021). When finetuning the model on CHILDES with 5 examples per episode, we use batch size (i.e., number of episodes/sequences per batch) $32$ and initial learning rate $3\times 10^{-3}$ and truncate the sequence to the max length of $80$ tokens to control the memory usage. When finetuning the model on CHILDES with 10 examples per episode, we use batch size $8$ and initial learning rate $3\times 10^{-4}$ and truncate the sequence to the max length of $180$ tokens. When finetuning the model on BabyLM-10M with $5$ examples per episode, we use batch size $16$ and initial learning rate $1\times 10^{-3}$ and truncate the sequence to the max length of $160$ tokens. Other settings are the same as when training from scratch except that we do not apply weight decay. Each run is performed on a single A100 GPU for 15 epochs on CHILDES (33 hours) or 12 epochs on BabyLM-10M (48 hours).
@@ Appendix C Held-out Word Classification
Here we provide an example task to further explain the classification paradigm in Section 3.3. Assume $K=4,~{}C=2$, and we reuse the example words and sentences in Figure 1. The word “ski” has its three study examples concatenated into a sequence:
@@
He will [new-word] past the pine trees. <sep>
@@
<sep> Look there’s an [new-word], it’s like an anteater <sep> See the [new-word] has a long snout for eating bugs. <sep> That must be the [new-word]’s house. <sep>
@@
The [new-word] is hungry, it wants some snacks. <sep>
When classifying the word “ski”, we compare the conditional likelihood of its query example “He will [new-word] past the pine trees. <sep>” in the following two sequences:
<sep> Susie learned to [new-token] last winter <sep> People [new-token] on tall mountains where there’s lots of snow <sep> I saw Susie [new-token] fast down the snowy mountain <sep> He will [new-word] past the pine trees. <sep>
<sep> Look there’s an [new-word], it’s like an anteater <sep> See the [new-word] has a long snout for eating bugs. <sep> That must be the [new-word]’s house. <sep> He will [new-word] past the pine trees. <sep>
@@
Below, we describe two baselines we run on this task.
Baseline: Llama-3 8B learning a pseudo-word in context (Llama-3 8B with ‘dax’)
@@
Additional Baseline: Llama-3 8B modeling the coherence of usage contexts (Llama-3 8B with ‘’)
@@
dataset $K$ $C$ Minnow from scratch Llama-3 8B with ‘’ Llama-3 8B with ‘dax Llama-3 8B +Minnow Llama-2 7B with ‘dax Llama-2 7B +Minnow Llama-2 7B +CoLLEGe
CHILDES $5$ 4 72.3(1.6) 58.33 71.09 79.1(0.5) 70.06 79.4(0.3) 62.45
@@
Table 6: Accuracy (%) of held-out word classification on the CHILDES and BabyLM-10M test sets. We show the mean and the standard deviation (in the bracket) of 3 runs. “Minnow from scratch” means models trained from scratch on the corresponding dataset. “Llama-3 8B with ‘”’ means the baseline model without prompt and remove the new word (i.e., replace the new word with an empty string). “Llama-3 8B with ‘dax”’ means the baseline model with prompt learning the new word ‘dax’. We use $K-1$ study examples in this classification task, and models except the baselines are trained/finetuned on $K$ examples per training episode so they see the same number of examples during training and evaluation. $C$ is the number of words in each group, so we will have $\lfloor\frac{n_{\text{episodes}}}{C}\rfloor$ groups. Note that we discard the last batch of less than $C$ episodes, so the used numbers of episodes are slightly smaller. Results of “Llama-3 8B with ‘”’ show that the coherence of the context already provides better-than-chance accuracy on this classification task. Results of “Llama-3 8B with ‘dax”’ show that the pre-trained LLM already performs well. However, “Llama-3 8B +Minnow” outperforms the baselines by a large margin, showing the effectiveness of our method. Models finetuned with Minnow from the instruction-tuned variant of Llama-3 8B perform worse than or close to the pre-trained variant here (the instruction-tuned variant finetuned with Minnow has 86.3% (4-way) and 80.1% (8-way) mean classification accuracies; the instruction-tuned variant with ‘dax’ has 75.2% (4-way) and 66.0% (8-way) classification accuracies), so we do not include their results here.
@@ Appendix D Syntactic Category Classification
Results from models trained from scratch, Llama-3 8B baseline and models finetuned from Llama-3 8B on the 6 category pairs and their mean are visualized in Figure 2. Table 7 shows detailed results from Llama-3 8B baseline and Llama-3 8B finetuned with Minnow on BabyLM-10M. Table 8 shows detailed results from models trained from scratch on both datasets.
@@
Table 7: LLMs’ accuracies (%) of distinguishing two syntactic categories in novel contexts. We show the mean and the standard deviation (in the bracket) of 3 runs. ‘Acc. (1$>$2)’ denotes the accuracy on the set of sentences where Category 1 should be preferred over Category 2 (e.g., assigning a higher probability to a noun in a noun-expecting context for row 1), and vice versa. Column ‘Acc.’ lists the aggregate accuracy. “Llama-3 8B +Minnow” have accuracies significantly better than chance except distinguishing adjective from verb (row 5). Additionally, “Llama-3 8B +Minnow” improves over Llama-3 8B baseline in differentiating most category pairs, showing the effectiveness of finetuning with Minnow.
@@
Table 8: Accuracies (%) of distinguishing two syntactic categories in novel contexts for models trained from scratch with Minnow. We show the mean and the standard deviation (in the bracket) of 3 runs. The formatting is the same as in Table 7. Both models perform better than chance on many category pairs, suggesting that models can develop some ability to one-shot learn the syntactic category of a word from human-scale data with Minnow.
Llama-2 7B +Minnow Llama-2 7B with CoLLEGe
Cat. 1 Cat. 2 Acc. Acc. (1$>$2) Acc. (2$>$1) Acc. Acc. (1$>$2) Acc. (2$>$1)
@@
Table 9: CoLLEGe’s Teehan et al. (2024) accuracies (%) of distinguishing two syntactic categories in novel contexts. The formatting is the same as in Table 7.
@@ Appendix E Comparing Generations
For new usage example generation (Section 5.3), we show GPT-4o the following text format:
@@
Tables 1 and 10 show the results of comparing Llama-3 8B baseline (pre-trained to instruction-tuned) to the Minnow model finetuned from that baseline (with random seed 0) on new examples and definitions generated by greedy decoding and top-p=$0.92$, respectively.
@@
Llama-3 8B Instruct baseline 46 52 33
+Minnow 47 36 28
Llama-2 7B +CoLLEGe 6 15 5
+Minnow 84 61 30
Table 10: Percentages of wins of each model when comparing the generations from Llama-3 8B baseline (pre-trained to instruction-tuned) with the Minnow model finetuned from that baseline, judged by GPT-4o. The left two datasets are for new usage example generation in Section 5.3, and the right-most one is for definition generation in Section 5.4. Each new example or definition is generated by top-p=$0.92$. The percentage of ties is the remaining after subtracting the win percentages of the two models. GPT-4o more frequently chooses the Minnow model as the winner compared to the corresponding baseline, except for the instruction-tuned model on Chimera and CoLLEGe-DefGen.

### v3

@@ Appendix A Word Usage Dataset Creation
Since the basic units of our focus are words (as opposed to word pieces in other tokenization schemes), we need to identify words in the text. To achieve this, we apply the same word-level tokenization to all datasets (for consistency) and mark word boundaries by whitespace during preprocessing. Models trained from scratch use this word-level tokenization. When the text is used in finetuning Llama, which comes with its pre-trained subword tokenizer, we remove the unnatural spaces introduced by the word-level tokenization and tokenize the text again with the Llama tokenizer, so the text format becomes closer to its pre-training data (See the Finetuning paragraph in Appendix B for further details of this process). For CHILDES data, we preprocess the data in the same way as Yedetore et al. (2023) did, which uses children’s input in the North American English portion, [^15] The version of CHILDES data we use is different from that of Yedetore et al. (2023), and the current version on the official webpage https://childes.talkbank.org/access/Eng-NA/ has also changed from our version. but we do not split and unk the data at the preprocessing stage. For BabyLM data, we use the data in the 10M track of the BabyLM Challenge 2023, which mixes 10 portions, each from a different data source (child- or adult-oriented, speech transcription or written text like Wikipedia). We exclude the QED portion for its poor quality (also mentioned in the 2nd BabyLM Challenge). We apply word-level tokenization on untokenized portions, and then split the text into sentences using heuristics. We use spaCy for all word-level tokenization along with Part-Of-Speech tagging. We lowercase all text before preprocessing to unify the capitalization of words in different places. We deduplicate sentences and remove sentences having less than 1 word (not counting punctuation).
@@ Appendix B Model and Training Configurations
We slightly modify the configuration of Pythia-160M (Biderman et al., 2023), which uses the Transformer architecture GPT-NeoX (Andonian et al., 2023). The configuration has $12$ layers and a hidden dimension size of $768$. We change the vocabulary size according to the corresponding dataset, as shown in Table 6. We also include three special tokens in the vocabulary: the placeholder token [new-token], the separator token <sep>, and <unk>, as mentioned in Section 4. We change the Pythia configuration to tie the input and output embeddings. This makes the model parameter counts smaller, 86.786.7M and 102.5102.5M for the model trained on CHILDES and BabyLM-10M, respectively. For both models, we use batch size (i.e., number of episodes/sequences per batch) $8$ and AdamW optimizer (Loshchilov and Hutter, 2019) with initial learning rate $3\times 10^{-4}$, and reduce the learning rate by multiplying $0.1$ when the validation loss has stopped improving for $2$ epochs. We apply weight decay $0.07$ and $0.15$ when training on the CHILDES and BabyLM-10M datasets, respectively. Other configurations, such as no dropout, are kept the same as Pythia-160M. For each setting, we run $3$ times with random seed $\{0,1,2\}$. Each run is performed on a single V100 GPU for 30 epochs (9–18 hours).
@@
We finetune Llama-3 8B (Meta AI, 2024) and Llama-2 7B (Touvron et al., 2023) with Minnow on each of the CHILDES and BabyLM-10M datasets, but we refer to the models finetuned on BabyLM-10M by default, as we mentioned in Section 5. We finetune from both the pre-trained and instruction-tuned variants of Llama-3 8B, but we refer to the models finetuned from the pre-trained variant by default, presenting results of finetuning from the instruction-tuned variant only in the generative settings, where their performance may differ considerably due to their different capabilities to follow the prompt. We finetune only the pre-trained variant of Llama-2 7B since that is what the CoLLEGe checkpoint is based on. We use two reserved special tokens in the Llama-3 tokenizer vocabulary (or two tokens added to the Llama-2 7B vocabulary) as the placeholder token and the separator token. To make the tokenization more natural to the model’s pre-training data, we clean up tokenization spaces in the text (e.g., the space before “,”, “.”, or “’s”) introduced by the word-level tokenization during preprocessing and make the placeholder token absorbs any preceding spaces of the word. Finetuning is minimally parameter-efficient: We finetune only the input and output embeddings of the two special tokens, while freezing all other parameters. Before finetuning, the input/output embedding of either token is initialized to the mean of all input/output embeddings (Hewitt, 2021). We finetune models on CHILDES with $5$ or $10$ examples per episode, and on BabyLM-10M with $5$ examples per episode. Detailed hyperparameters we use to finetune Llama-3 8B and Llama-2 7B are summarized in Table 7. Other settings are the same as when training from scratch except that we do not apply weight decay. Each run is performed on a single A100 GPU for 15 epochs on CHILDES (33 hours) or 12 epochs on BabyLM-10M (48 hours).
dataset $K$
batch
size
max. seq.
length
initial learning rate
Llama-3 8B Llama-2 7B
CHILDES 5 32 80 $3\times 10^{-3}$ $1\times 10^{-3}$
10 8 160 $3\times 10^{-4}$ $1\times 10^{-3}$
BabyLM-10M 5 16 160 $1\times 10^{-3}$ $3\times 10^{-3}$
Table 7: Finetuning hyperparameters for different datasets and settings. $K$ is the number of examples per episode. “batch size” is the number of episodes/sequences per batch. “max. seq. length” is the maximum number of tokens we truncate the sequence to in order to control the memory usage.
@@ Appendix C Held-out Word Classification
Here we provide a full explanation of the example task in Figure 2 to further explain the classification paradigm in Section 3.3. Assume K=4,C=2K=4,~C=2, and we reuse the example words and sentences in Figure 1. As Figure 2 shows, the word “ski” has its three study examples concatenated into a sequence:
@@
He will [new-token] past the pine trees. <sep>
@@
<sep> Look there’s an [new-token], it’s like an anteater <sep> See the [new-token] has a long snout for eating bugs. <sep> That must be the [new-token]’s house. <sep>
@@
The [new-token] is hungry, it wants some snacks. <sep>
When classifying the word “ski”, we compare the conditional likelihood of its query example “He will [new-token] past the pine trees. <sep>” in the following two sequences:
<sep> Susie learned to [new-token] last winter <sep> People [new-token] on tall mountains where there’s lots of snow <sep> I saw Susie [new-token] fast down the snowy mountain <sep> He will [new-token] past the pine trees. <sep>
<sep> Look there’s an [new-token], it’s like an anteater <sep> See the [new-token] has a long snout for eating bugs. <sep> That must be the [new-token]’s house. <sep> He will [new-token] past the pine trees. <sep>
@@
Below, we describe two kinds of baselines we run on this task.
Baseline: pre-trained LLM learning a pseudo-word in context (Llama-3 8B or Llama-2 7B with ‘dax’)
@@
Additional Baseline: pre-trained LLM modeling the coherence of usage contexts (Llama-3 8B with ‘’)
@@
dataset $K$ $C$
Minnow
from scratch
Llama-3 8B
with ‘’
Llama-3 8B
with ‘dax
Llama-3 8B
+Minnow
Llama-2 7B
with ‘dax
Llama-2 7B
+Minnow
Llama-2 7B
+CoLLEGe
CHILDES 55 4 72.3(1.6) 58.33 71.09 79.1(0.5) 70.06 79.4(0.3) 62.45
@@
Table 8: Accuracy (%) of held-out word classification on the CHILDES and BabyLM-10M test sets. We show the mean and the standard deviation (in brackets) of 3 runs. “Minnow from scratch” means models trained from scratch on the corresponding dataset. “Llama-3 8B with ‘”’ means the baseline model without prompt and remove the new word (i.e., replace the new word with an empty string). “Llama-3 8B with ‘dax”’ or “Llama-2 7B with ‘dax”’ means the baseline model with prompt learning the new word ‘dax’. We use K1K-1 study examples in this classification task, and models except the baselines are trained/finetuned on $K$ examples per training episode so they see the same number of examples during training and evaluation. $C$ is the number of words in each group, so we will have $\lfloor\frac{n_{\text{episodes}}}{C}\rfloor$ groups. Note that we discard the last batch of less than $C$ episodes, so the used numbers of episodes are slightly smaller. Results of “Llama-3 8B with ‘”’ show that the coherence of the context already provides better-than-chance accuracy on this classification task. Results of “Llama-3 8B with ‘dax”’ show that the pre-trained LLM already performs well. However, “Llama-3 8B +Minnow” outperforms the baselines by a large margin, showing the effectiveness of our method. Models finetuned with Minnow from the instruction-tuned variant of Llama-3 8B perform worse than or close to the pre-trained variant here (the instruction-tuned variant finetuned with Minnow has 86.3% (4-way) and 80.1% (8-way) mean classification accuracies; the instruction-tuned variant with ‘dax’ has 75.2% (4-way) and 66.0% (8-way) classification accuracies), so we do not include their results here. Similar improvements by Minnow is also shown on Llama-2 7B by comparing the baseline “Llama-2 8B with ‘dax”’ and “Llama-2 8B +Minnow”. Moreover, the CoLLEGe baseline (“Llama-2 8B +CoLLEGe”) performs even worse than the baseline.
@@ Appendix D Syntactic Category Classification
Results from models trained from scratch, Llama-3 8B and Llama-2 7B baseline, models finetuned from Llama-3 8B and Llama-2 7B, and the CoLLEGe baseline on the 6 category pairs and their mean are visualized in Figure 3. Table 9 shows detailed results from Llama-3 8B baseline and Llama-3 8B finetuned with Minnow on BabyLM-10M. Table 10 shows detailed results from models trained from scratch on both datasets. Table 11 shows detailed results from Llama-2 7B finetuned with Minnow on BabyLM-10M and the CoLLEGe baseline.
@@
Table 9: Comparing Llama-3 8B baseline and the Minnow model finetuned from it on their accuracies (%) of distinguishing two syntactic categories in novel contexts. We show the mean and the standard deviation (in brackets) of 3 runs. Following Table 1 in Kim and Smolensky (2021), ‘Acc. (1>>2)’ denotes the accuracy on the set of query sentences where Category 1 should be preferred over Category 2 (e.g., for row 1, assigning a higher likelihood to the noun-expecting query sentence when the placeholder represents a noun compared to a verb; using the examples in Section 5.2, the query example (1) expecting $w^{(1)}$ is among this set of query sentences, and it should have a higher likelihood when [new-token] represents $w^{(1)}$ compared to $w^{(2)}$.), and vice versa. Column ‘Acc.’ lists the aggregate accuracy. “Llama-3 8B baseline” generally has accuracies better than chance (50%) except for distinguishing certain pairs of categories. Additionally, “Llama-3 8B +Minnow” improves over Llama-3 8B baseline in differentiating most category pairs, showing the effectiveness of finetuning with Minnow.
@@
Table 10: Accuracies (%) of distinguishing two syntactic categories in novel contexts for models trained from scratch with Minnow. We show the mean and the standard deviation (in brackets) of 3 runs. The formatting is the same as in Table 9. Both models perform better than chance on many category pairs, suggesting that models can develop some ability to one-shot learn the syntactic category of a word from human-scale data with Minnow. In general, models trained on BabyLM-10M perform better than models trained on CHILDES, probably because the BabyLM dataset is more diverse and contains formal written texts, closer to the MNLI dataset, from which this test dataset is built.
Llama-2 7B +Minnow Llama-2 7B +CoLLEGe
Cat. 1 Cat. 2 Acc. Acc. (1>>2) Acc. (2>>1) Acc. Acc. (1>>2) Acc. (2>>1)
@@
Table 11: Comparing the Minnow models finetuned from Llama-2 7B and CoLLEGe Teehan et al. (2024) on their accuracies (%) of distinguishing two syntactic categories in novel contexts. The formatting is the same as in Table 9. Overall, both models perform well on the task (they both have 80% mean accuracy as mentioned in Figure 3). The CoLLEGe model performs better in more settings, but it fails to distinguish verbs from nouns (row 1, last column).
@@ Appendix E Comparing Generations
As we mentioned in Sections 5.3 and 5.4, for the quantitative evaluation, we compare a pair of generations (new usage examples or definitions) from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. In addition to GPT-4o evaluation, we have also asked the first author to conduct a small-scale human evaluation to compare Llama-3 8B baseline and a Minnow model finetuned from it. [^16] We acknowledge this is very limited and have problems, and leave larger-scale systematic human evaluation for future work. We show GPT-4o and the human the same prompts.
For new usage example generation (Section 5.3), we show GPT-4o or human the following text format:
@@
GPT-4o evaluation
Tables 3 and 12 show the results of comparing pairs of new usage examples or definitions generated from Llama-3 8B baseline (pre-trained to instruction-tuned) to a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B, by greedy decoding and top-p=$0.92$, respectively.
@@
Llama-3 8B Instruct baseline 43 46 33
+Minnow 46 38 29
Llama-2 7B +CoLLEGe 5 11 5
+Minnow 85 68 30
Table 12: Percentages of wins of each model when comparing the generations from the pairs of models in each box, judged by GPT-4o. In the top two ruled rows, we compare Llama-3 8B baseline (pre-trained to instruction-tuned) with a Minnow model finetuned from that baseline (averaged across 3 runs). In the bottom-most ruled row, we compare CoLLEGe with the Minnow model finetuned from the pre-trained Llama-2 7B (averaged across 3 runs). The left two datasets are for new usage example generation (Section 5.3; each new usage example is generated by providing 4 study examples), and the right-most one is for definition generation (Section 5.4; each definition is generated by providing 3 study examples). Each new example or definition is generated by top-p=$0.92$. We boldface significantly more preferred models ($p<.05$ in paired t-tests across 3 runs). The percentage of ties is the remaining after subtracting the win percentages of the two models. GPT-4o prefers the Minnow model compared to the pre-trained Llama-3 8B baseline on two datasets. GPT-4o also strongly prefers the Minnow model compared to the CoLLEGe baseline ($p=0.01$ on Chimera and $p<.001$ on other two datasets). The difference made by Minnow on the instruction-tuned Llama-3 8B are not significant ($p>.1$).
Human evaluation
Tables 13 and 14 show the results of comparing pairs of new usage examples or definitions generated from Llama-3 8B baseline (pre-trained) to a Minnow model finetuned from it, by greedy decoding and top-p=$0.92$, respectively.
New Example Definition
Variant Method BabyLM-10M test Chimera CoLLEGe-DefGen
Llama-3 8B baseline 6 9 20
+Minnow 32 30 22
Table 13: Percentages of wins of each model when comparing the generations from Llama-3 8B baseline (pre-trained) with a Minnow model finetuned from that baseline (with random seed 0), judged by the human. Each new example is generated by greedy decoding. The percentage of ties is the remaining after subtracting the win percentages of the two models. Due to the high cost of human evaluation, the human evaluates 50 pairs sampled from each of the BabyLM-10M test portion and the CoLLEGe-DefGen dataset. The human more frequently choose the Minnow model as the winner compared to the baseline.
New Example Definition
Variant Method BabyLM-10M test Chimera CoLLEGe-DefGen
Llama-3 8B baseline 18 24 22
+Minnow 38 27 28
Table 14: Percentages of wins of each model when comparing the generations from Llama-3 8B baseline (pre-trained) with a Minnow model finetuned from that baseline (with random seed 0), judged by the human. Each new example is generated by sampling with top-p=$0.92$. The percentage of ties is the remaining after subtracting the win percentages of the two models. Due to the high cost of human evaluation, the human evaluates 50 pairs sampled from each of the BabyLM-10M test portion and the CoLLEGe-DefGen dataset. The human more frequently choose the Minnow model as the winner compared to the baseline.
To examine how the human and GPT-4o agree and disagree, Tables 1516 and 17 show the counts of generations with each pair of human-GPT-4o judgments on the BabyLM-10M test portion, the Chimera dataset, and the CoLLEGe-DefGen dataset, respectively. In general, GPT-4o still underestimates the improvement of the Minnow model compared to the baseline, enhancing the conclusion regarding the effectiveness of our method.
human GPT-4o +Minnow baseline tie
+Minnow 26 7 2
baseline 3 9 0
tie 20 23 10
Table 15: Comparison between judgments made by GPT-4o and the human on the BabyLM-10M test portion. The human evaluates 50 words, each has two pairs of generations: one generated by greedy decoding and one generated by top-p=$0.92$ sampling, resulting in 100 pairs in total. By comparing the off-diagonal numbers, we know that when the human and GPT-4o disagree, GPT-4o tends to favor the baseline, which suggests that GPT-4o still underestimates the improvement brought by finetuning with Minnow compared to the human.
human GPT-4o +Minnow baseline tie
+Minnow 17 0 2
baseline 1 8 2
tie 14 19 3
Table 16: Comparison between judgments made by GPT-4o and the human on the Chimera dataset. There are 33 chimera words, each has two pairs of generations: one generated by greedy decoding and one generated by top-p=$0.92$ sampling, resulting in 66 pairs in total.
human GPT-4o +Minnow baseline tie
+Minnow 15 4 6
baseline 3 15 3
tie 8 14 32
Table 17: Comparison between judgments made by GPT-4o and the human on the CoLLEGe-DefGen dataset. The human evaluates 50 words, each has two pairs of generations: one generated by greedy decoding and one generated by top-p=$0.92$ sampling, resulting in 100 pairs in total.